# Upgrading Democracies with Fairer Voting Methods

### v1

@@ 2.1 Fair ballot aggregation has distinguishing multi-faceted qualities
Figure 2: Compared to earlier participatory budgeting experience based on the standard utilitarian greedy method, the method of equal shares increases the representation of voters’ preferences and project categories, in both project ideation and selection phase. (A) The representation of voter preferences increases by 14.5% under equal shares compared to earlier participatory budgeting processes based on the standard utilitarian greedy method. (B) The projects proposed for voting under equal shares promote project categories that have traditionally been under-represented, such as welfare and health, while projects that usually come with high costs, such as public transit, lose appeal significantly. (C) It is also these otherwise under-represented projects that become the winning ones when using equal shares. On average, project categories show a 3% higher representation.
@@
Equal shares encourages proposed projects that are novel and win. Figure 2b illustrates the share of the proposed project categories of participatory budgeting in Aarau compared to the projects observed in earlier participatory budgeting processes [33, 32]. We find a significant change in the nature of the proposed projects: welfare and health projects, which were usually under-represented, are more prominently featured, while public transit projects are less prominent. This shift is determined by the costs of the projects belonging in the two categories: the average cost of public transit projects in the participatory budgeting process of Aarau amounts to 8% of the overall budget, while welfare-related projects amount to 12.25%. In earlier participatory budgeting processes, these numbers were 18.19% and 7.52%. Figure 2c presents the share of the winning project categories in the participatory budgeting process of Aarau for the method of equal shares and for the majoritarian standard method. It also shows the share in earlier participatory budgeting experience that relied on the standard method. Health and sport would be disadvantaged with the standard method as it would reduce their share in the winning set by 19.13% and 18.05%, respectively.
@@
Broader evidence beyond the case of Aarau. To strengthen further our findings, we also highlight the case of two other participatory budgeting elections that used equal shares during 2023 in two cities of Poland: (i) Świecie and (ii) Wieliczka (“"Green Million"). In the first case that was a follow up of Aarau, 17/22 projects are selected for funding, while the majoritarian standard method would result in only 9 winning projects. The representation of voters’ preferences increased from 61% to 80%,. The top-2 project in voters’ approvals is not funded, which requires 36.6% of the budget and is approved by 24.4% of the population. In the second case, 30/64 projects are selected for funding, while the majoritarian standard method would result in 23 winning projects. The representation of voters preferences increased from 61% to 71%,. The top-2 project in voters’ approvals is not funded, which requires 10% of the budget and is approved by 8.9% of the population. Both cases confirm the distinguishing voting outcomes of equal shares: proportional representation is achieved by funding more projects that increase the overall satisfaction with compromises on popular expensive projects. In contrast to these two elections, further novel data are collected during the participatory budgeting process of Aarau that allow us to rigorously explore the legitimacy of fair voting methods and which citizens’ values explain the shift to those.
2.2 Overcoming the adoption barrier of fair voting improves legitimacy
@@
Adopting a new ballot aggregation method: a rigorous evaluation. Figure 3 illustrates which ballot aggregation method voters prefer and which they find fairer in the entry survey before voting and in the exit survey after voting. Section LABEL:sec:choice-ballot-aggregation also shows additional results of which method voters find more convenient. Section LABEL:sec:equal-shares-explanation further illustrates how the two methods are presented to voters, in particular the voting outcomes and the explanations, based on which the measurements are made. The shift from the majoritarian standard method to the method of equal shares becomes obvious by the gain in the share of voters, who prefer and find fairer equal shares after voting (rigorously measured on the same matched population). We also assessed which method voters found more convenient in Figure LABEL:fig:convenience. The legitimacy of the two ballot aggregation methods is determined by measuring the satisfaction of voters with the voting outcomes of each method (even if the voter classifies as a ‘loser’, based on different measurements), see Section 4.2 for more information. To assess legitimacy thoroughly, the voters’ gain is disaggregated further down to all 24 combinations of winners/losers and satisfied/unsatisfied voters (see Figure LABEL:fig:legitimacy_fairness_preference_heat). We group those into three groups in Figure 3 for simplicity: satisfied winners, unsatisfied losers and mixed, each consisting of 8 combinations. Three ways to measure each of the winners/losers and satisfied/unsatisfied are illustrated in Section 4.2.
@@
Figure 3: In the exit survey after voting, the method of equal shares is found more preferred and fairer than the majoritarian standard ballot aggregation method. Strikingly, before voting and adopting equal shares, there is no clear preference for any of the two methods, while the standard method is found fairer. Voting outcomes themselves are sufficient for individuals to determine equal shares as their preferred and fair ballot aggregation method. The explanation of the methods does not significantly contribute to shifting the choice to equal shares. The legitimacy of equal shares is assessed based on voters’ (A) preference and perceived (B) fairness of equal shares over the standard method. The x-axis determines the difference of the voters share between those choosing equal shares over the standard method. Legitimacy is assessed based on 24 combinations of winners/losers and satisfied/unsatisfied voters that are grouped into (i) winners and satisfied, (ii) losers and unsatisfied and (iii) mixed. Winners/losers are determined in three different ways: 1st and 2nd quartile of representation score for the winning projects that are (i) non-exclusive and (ii) exclusive between the two ballot aggregation methods. (iii) The non-supporters and supporters of the top popular losing project. Satisfied/unsatisfied voters are determined by three post-voting survey questions that assess: (i) satisfaction, (ii) acceptance of outcome and (iii) accurate representation of will.
@@
The adoption of equal shares exhibits a shift to be the most preferred and fairer method. The results of Figure 3 show that, in the exit survey after voting as compared to the entry survey before, there is a significant gain in voters who prefer the method of equal share and find it fairer. This is particularly prominent for unsatisfied losers who choose the standard method as fairer before voting but switch to equal shares after voting (see Figure LABEL:fig:legitimacy_fairness_preference_heat). More specifically, 54% of the population prefers equal shares before voting, while after voting, supporters of equal shares raise to 80% (p=0.0034). Similarly, 45% of voters find equal shares fairer before voting and 89% fairer after voting (p=0.028).
@@
The method of equal shares shows stronger legitimacy. The significant legitimacy of equal shares is reflected in several ways (see also Figure LABEL:fig:legitimacy_fairness_preference_heat): (i) Losers, determined as supporters of the top popular losing project (‘Wild bees’ paradise’), prefer the method of equal shares and find it fairer. (ii) Remarkably, the method of equal shares is preferred and found fairer even more among losers than among winning voters, who prefer the majoritarian standard method. (iii) The shift to equal shares after voting is even found among those who do not accept the voting outcome, which is a strong proxy of legitimacy [34, 35].
When shift to equal shares is not achieved. However, we also found some cases where more voters shifted to the majoritarian standard method after voting, namely (i) satisfied voters preferring the standard method over unsatisfied voters preferring equal shares; (ii) voters accepting the voting outcome and finding the standard method fairer over voters who do not accept the voting outcome and find equal shares fairer; (iii) unsatisfied losers (or winners), who prefer the standard method over unsatisfied winners (or losers), who choose equal shares, when winners/losers are determined by the representation and their choice for the top popular losing project (‘Wild bees’ paradise’). See Figure LABEL:fig:legitimacy_fairness_preference_heat for a more disaggregated view.
@@
Voting outcome itself vs. method explanation: what matters. The method explanation has an additional contribution of 5.3% to the shift to equal shares as preferred method, while it has an addition contribution of 5.2% to the shift to equal share as fairer method. The method explanation has a small but statistically significant contribution to the shift to equal shares (with only exception the mixed setting for the preferred method, p=0.07). These results demonstrate that the voting outcome itself is sufficient to establish a change.
@@
2.3 Democratic values explain voters’ shift to fairer voting methods
@@
Towards fairer voting: A clash between democratic values and democratic backsliding. Figure 4 identifies distinguishing human factors that causally explain the shift from one ballot aggregation method before voting to another one after voting (with the explanation of method), as illustrated in Figure 3 and in Section 4.3. The classification model is illustrated in Section LABEL:sec:shifts-methods and comes with an accuracy of 85.92%. Strikingly, the group of voters who shift from the majoritarian standard method to equal shares exhibits democratic values that explain this transition, see Figure 4c. These voters come with distinguishing qualities (Figure 4b) such as being altruistic, compromisers and supporting more expressive, inclusive ballots. They are aware of the proposed projects and flexible to select low-cost projects, which also makes them more likely to oppose a popular costly project. This group is well-represented in the voting outcome. In contrast, the voters that shift from equal shares to the majoritarian standard method do not exhibit these democratic values, but are driven by self-interest, mistrust and dissatisfaction. They support simpler ballot formats, are not well-represented in the voting outcomes, and support the popular loser project under equal shares.
@@
Figure 4: The shift to equal shares as the preferred method after voting is explained by distinguishing human traits that reflect the promotion of strong democratic values such as being altruistic, compromisers and supporting more expressive, inclusive ballots. In contrast, shifting to the majoritarian standard method is found among voters showing mistrust, under-representation, dissatisfaction, and self-interest. These voters also support the popular costly loser project of the equal shares and simple ballot methods. A classification model (with 85.92% accuracy) is designed with four classes representing the preferred ballot aggregation method before and after voting (with explanation of methods). The independent variables are human traits of voters that explain the observed transitions. (A) Distinguishing human traits of voters who shift to the standard method and equal shares after voting. These are prominently associated with democratic backsliding on the one hand, and promotion of democratic values on the other hand. (B) Summary of distinguishing qualities for those who prefer the method of equal shares over the majoritarian standard method after voting (feature difference of the two classes). Altruism, support of expressive, inclusive ballots and opposition to the popular, costly loser project of equal shares are qualities with the highest prominence for those voters preferring equal shares. (C) Human traits (features) that explain the preference of each voter before and after voting, in particular the transitions from the majoritarian standard method to equal shares and from equal shares to the standard method (see Figure 3). The ‘*’ indicates that a feature is statistically significant (p<0.05) in explaining the voters’ choice of the method before and after voting.
@@
Distinguishing qualities of voters who shift preference to equal share. Figure 4b summarizes the distinguishing qualities of voters, who shift to equal shares by taking the difference of the human traits for the two transitions (standard to equal shares - equal shares to standard). Altruism, opposition to the popular (costly) loser project and support of the expressive, inclusive ballot formats are the most prominent distinguishing qualities observed for voters who shift to equal shares. Figure 4c shows the importance of features that explain the classification of voters based on which of the two transitions they make for their preferred ballot aggregation method before and after voting (see Tables LABEL:tab:accuracy_resultsLABEL:tab:feature-loss and LABEL:tab:trait-survey-map in Section LABEL:sec:shifts-methods). Further results for the transitions to methods that are found fairer and more complex are shown in Figures LABEL:fig:ES_G_G_ESLABEL:fig:ES_GLABEL:fig:Convenience_Explain in Section LABEL:sec:shifts-methods.
@@ 3 Discussion
A blueprint to upgrade democracies. In our study, we have shown how voting can better foster collective intelligence that is based on higher inclusion and diversity, with strong evidence provided by a rigorous participatory budgeting process in the city of Aarau in Switzerland. Subsequent participatory budgeting processes run using this approach further confirm our findings [36]. In particular, we found that the majoritarian standard approach of voting had a number of deficits, such as the selection of a few projects, leaving a large portion of the project ideas unfulfilled. As the selected projects are implemented mainly in the city center, it largely fails to reach spatial fairness or inclusion. However, combining cumulative voting with the method of equal shares proves to be resourceful, as it promotes more novel winning projects, fairness, satisfaction, inclusion, trust and democratic values. Accordingly, this approach offers a powerful way of upgrading democracies. Implementing this approach, using digital means, promises to be a powerful, value-sensitive approach to promote digital democracy by design [7].
@@
Success of project implementation. By the end of 2024, 9 of 17 projects selected by the "City Idea" process in Aarau were already implemented and 5 projects got an extension, while 3 projects could not be implemented mainly due to feasibility and budgetary constraints [29]. These were among the top-11 projects. Two thirds of these projects would also have been winners if the majoritarian standard method had been used. These results demonstrate that having more winning projects using the equal shares method does not negatively affect the overall capacity of citizens to implement them.
@@
Participation levels. The participatory budgeting process in Aarau achieved significant participation of citizens at different levels. During the project ideation phase, 161 ideas were proposed, which exceeded the initial goal of the campaign to attract 100 ideas by far [29]. The total of 1703 voters who participated is a significant number for such a campaign in a town with 14,310 residents having voting rights in 2023. Nevertheless, the turnout did not reach the level of official elections such as referenda and national elections. A 22% of the voters showed interest in participating, although they were not officially residents of Aarau, while 9% of the voters were not Swiss citizens. The average age of voters was 43, and there were more female than male voters. Children over 15 years old were allowed to vote. Among the low number of voters who abstained but participated in the surveys (296), the top-5 reasons for not voting were "no time" (17.1%), "forgetting" (12.8%), "lack of information" (9.9%), "unappealing projects" (7.7%) or some "other reason" (7.0%), see Question EB.7 in Table LABEL:table:exit-digital.
@@
Upgrading democracies requires upgrading digital tools. We created and used a novel combination of open-source and value-sensitive digital tools to meet a broad range of expectations regarding transparency, security, privacy, reproducibility, explainability and fairness. This required a significant implementation effort, a mobilization of various communities, a strong commitment and the building of shared values with the city authorities, which is a process that has not been systematic pursued to this extent so far. A highlight of this effort was the privacy-preserving verification of citizens’ eligibility when voting with the social security number, which ensured a trusted voting procedure. The integration of our open-source implementation of cumulative voting [30] into the Stanford Participatory Budgeting platform is a stepping stone towards a more ambitious roadmap, which aims to make participatory tools more standardized and interoperable, in particular the ones supported by large open-source communities such as Decidim and Consul as well as the ones created by the computational social choice community such as Pabutools [37] and Preflib [38].
@@
Explaining and adopting fair voting methods. A 53.17% of voters understood how equal shares calculates the winning projects (MB.3.3), while a 55.95% understood why popular projects are not easily funded by equal shares (MB.3.6), and 41.26% find that fair (MB.3.7). Strikingly, the results of Figure 3 demonstrate that explaining the expected difference in the voting outcomes reached by equal shares is sufficient for voters to prefer this method. A more detailed explanation of how the method works increases the tendency towards preferring equal shares over the standard method only slightly. This result has significant implications for overcoming the long-standing hesitation with regard to democratic innovations, since citizens show a readiness to accept more complex decision-making methods as long as they understand the anticipated impact on the decision outcome. In addition, it helps when people experience the different result, which gives them a practical understanding of the method. Moreover, given that equal shares is designed for a proportional representation of voter preferences and is known for its stability [19], it is more unlikely to yield controversies over voting outcomes. Nevertheless, the additional effort required to explain the method of equal shares should not be under-estimated. As shown in Section LABEL:sec:equal-shares-explanation, the introduction of equal shares requires a systematic communication effort over the lifetime of the participatory budgeting process, using creative media and visualizations suited for the broader public. At the project ideation phase, awareness of the method is required to mobilize different parts of the populations for ideas that may not have been considered under the majoritarian rule. As the costs of the proposed projects play a key role for selecting them under proportional representation, citizens need to be resourceful and prioritize different ideas, meaning a focus on cost-effective projects that get the support by a sufficiently large group of the population. Large and costly projects may be more risky to propose as they require the support of a broader population to get selected.

### v2

@@ 2.2 Generalizing fair voting methods as a distinguishing paradigm
Figure 2: Compared to earlier participatory budgeting experiences based on the standard utilitarian greedy method, the method of equal shares improves the representation of voter preferences and project categories, both in the project ideation and the project selection phase. (A) 27.8 % more voters on average get any minimum representation level using the method of equal shares as compared to earlier participatory budgeting processes based on the standard utilitarian greedy method. (B) The projects proposed for voting under equal shares promote project categories that have traditionally been under-represented, such as welfare and health, while projects that usually come with high costs, such as public transit, lose appeal. (C) When using the method of equal shares, some otherwise under-represented projects are winning. On average, project categories show a 3% higher representation.
Strikingly, much more voters (difference of 57.7% to 80.7%) are well-represented to a minimum extent with equal shares in Aarau, even when considering up to the minimum representation of 70% (Figure 2a). In contrast, a larger voters share for more than 90% representation is observed for the past elections using the standard majoritarian method. Such so high representation levels imply less capacity for compromise, leaving a larger part of the population with lower representation levels. This representation gain is not so high (difference of 49.6% to 56.7%) if the method of equal shares is applied to the past 39 past elections. This can be explained by (i) the different nature of the proposed projects when using equal shares, coming with lower cost (48.4% lower budget share) and (ii) the larger number of winning projects (23.8% relative increase in the number of winning projects). Table LABEL:tab:representation-gain also confirms that participatory budgeting processes with the project and budget characteristics (matching criteria) of Aarau result in the highest representation gain. This strengthens further the evidence for the approach of Aarau as a blueprint for democratic updates based on fair voting methods.
@@
The method of equal shares encourages proposed projects that are novel and win. Figure 2b illustrates the share of the proposed project categories in the participatory budgeting in Aarau compared to projects proposed in earlier participatory budgeting processes [Faliszewski2023, pabulib]. We find a significant change in the nature of the proposed projects: welfare and health projects, which were usually under-represented before, are more prominently featured, while public transit projects are less prominent. This shift is reflected by the costs of projects in two categories: the average costs of public transit projects in the participatory budgeting process of Aarau amount to 8% of the overall budget, while welfare-related projects amount to 12.25%. In earlier participatory budgeting processes, these numbers were 18.19% and 7.52%, respectively. Figure 2c presents the share of the winning project categories in the participatory budgeting process of Aarau for the method of equal shares and for the standard majoritarian method. It also shows the share in earlier participatory budgeting experiences that relied on the standard method. Health and sports would be disadvantaged by the standard method, as it would reduce their share in the winning set by 19.13% and 18.05%, respectively.
Regional comparison: Aarau [equal shares] vs. Zurich [standard]. To further demonstrate the impact and validity of the innovative voting approach used in Aarau, we compare it with the previous participatory budgeting process in Zurich, which used the standard majoritarian voting method. Zurich (which is 38 km away from Aarau, i.e. located in rather close proximity) ran the “City Idea” format in 2021 [ZurichPBWeb, ZurichPBReport]. Although data of individual voters are not available for analysis to us, a comparison based on aggregate measures is a particularly revealing ‘crash-test’ because of the following interesting facts:
  • (i)
    With a population size of 423,193>22,013, a larger pool of citizens can be attracted to participate in the “City Idea” process in Zurich as compared to Aarau;
  • (ii)
    The winning/proposed projects originated from all quarters of Zurich, which aimed to encourage a broad engagement by the citizens of Zurich in the participatory budgeting process: West: 16/33, North: 14/24, East: 14/29, South: 17/49;
  • (iii)
    The available budget for all regions was around 540,000 CHF, which was much more than the 55,000 CHF provided by the smaller city of Aarau.
Despite those distinguishing factors, which are in favor of Zurich, the approach in Aarau turned out to be more effective than the one in Zurich. Even though Zurich had a larger size and a more diverse geographical origin of its proposed projects, the process attracted only the votes of 1,804 citizens, which was approximately the same number as in Aarau, corresponding to a 17 times higher turnout in Aarau. Moreover, many of the proposed projects in Zurich claimed the maximum amount of 10,000 CHF, while the average cost was 6.6% of the total budget, which is even lower than the 15% in Aarau.
Accordingly, the projects in Aarau came with different cost modalities and higher diversity. Although the standard majoritarian voting method used in Zurich prioritized popularity, there were 7 projects in Zurich that were not selected because they each fell short by just 1 to 4 additional votes. This creates several origins of dissatisfaction for the voters in Zurich, in contrast to the process in Aarau, where the fair voting methods create several new ways for voters to have their preferred projects winning as showed in Section 2.1.
@@
Broader evidence beyond the case of Aarau. To strengthen our findings further, we also highlight the case of two other participatory budgeting processes that used the method of equal shares during 2023 in two cities of Poland: (i) Świecie and (ii) Wieliczka (“Green Million”). In the first case of Świecie that was a follow-up of Aarau, 17 of 22 projects are selected for funding, while the majoritarian standard method would have resulted in only 9 winning projects. In addition, the representation of voters’ preferences increased from 61% to 80% by the method of equal shares. The top-2 project in terms of voter approvals was not funded, which would have required 36.6% of the budget, while it was approved by 24.4% of the population. In the second case of Wieliczka, 30 of 64 projects were selected for funding, while the majoritarian standard method would have resulted in 23 winning projects. Here, the representation of voters’ preferences increased from 61% to 71%. The top-2 project in voters’ approvals was also not funded, which would have required 10% of the budget and was approved by 8.9% of the population. Both cases confirm the distinguishing voting outcomes we found for the method of equal shares: proportional representation is achieved by funding more projects, while a few expensive popular projects are sacrificed. However, this increases the overall satisfaction. In contrast to these two elections, further novel data were collected during the participatory budgeting process in Aarau, which allow us to rigorously explore the legitimacy of fair voting methods and the citizen values that explain the shift to those.
2.3 Adopting fair voting improves legitimacy
@@
Rigorous evaluation of a new ballot aggregation method. Figure 3 illustrates which ballot aggregation method voters prefer and which they find fairer in the entry survey before voting and in the exit survey after voting. Section LABEL:sec:choice-ballot-aggregation also shows additional results of which method voters find more convenient. Section LABEL:sec:equal-shares-explanation further illustrates how the two methods, based on which our measurements were made, were presented to the voters, in particular the voting outcomes and the explanations. The shift from the majoritarian standard method to the method of equal shares becomes obvious by the gain in the share of voters, who prefer equal shares after voting and find it fairer (rigorously measured on the same matched population). We also assessed which method voters found more convenient (see Figure LABEL:fig:convenience). The legitimacy of the two ballot aggregation methods was determined by measuring the satisfaction of voters with the voting outcomes of each method. Remarkably, it increased even if the voter is classified as a ‘loser’ (based on different measurements); see Section 4.2 for more information. To assess legitimacy thoroughly, the voters’ gain (from before to after voting) was disaggregated further down to all 24 combinations of winners/losers and satisfied/unsatisfied voters (see Figure LABEL:fig:legitimacy_fairness_preference_heat). For simplicity, in Figure 3 we grouped those into three groups: satisfied winners, unsatisfied losers and mixed, each consisting of 8 combinations. Three ways to measure each of the winners/losers and satisfied/unsatisfied are illustrated in Section 4.2.
@@
Figure 3: In the exit survey after voting, the method of equal shares was found to be more preferred and fairer than the standard majoritarian ballot aggregation method. Strikingly, before voting and adopting equal shares, there was no clear preference for any of the two methods, while the standard method was found fairer. The voting outcomes themselves were sufficient for individuals to determine equal shares as their preferred and fairer ballot aggregation method. (The explanation of the methods does not significantly contribute to shifting the choice to equal shares.) Furthermore, the legitimacy of equal shares was assessed based on (A) the voters’ preference and (B) their perceived fairness of equal shares over the standard method. The x-axis determines the difference (gain) of the voters proportion between those choosing equal shares and those choosing the standard method. Legitimacy is assessed based on 24 combinations of winners/losers and satisfied/unsatisfied voters that are grouped into (i) satisfied winners, (ii) unsatisfied losers and (iii) mixed. Winners/losers are determined in three different ways: 1st and 2nd quartile of representation score for the winning projects that are (i) non-exclusive and (ii) exclusive between the two ballot aggregation methods. (iii) Non-supporters and supporters of the top popular losing project. Satisfied/unsatisfied voters are determined by three post-voting survey questions that assess: (i) satisfaction, (ii) acceptance of the outcome and (iii) accurate representation of will.
@@
The adoption of the method of equal shares exhibits a shift to be the most preferred and fairer method. The results of Figure 3 show that, there is a significant gain in voters who prefer the method of equal share and find it fairer in the exit survey after voting when compared to the entry survey before. This is particularly prominent for unsatisfied losers who chose the standard method as fairer before voting but switched to equal shares after voting (see Figure LABEL:fig:legitimacy_fairness_preference_heat). More specifically, 54% of the population preferred equal shares before voting, while after voting, supporters of equal shares raised to 80% (p=0.0034). Similarly, 45% of voters found equal shares fairer before voting and 89% fairer after voting (p=0.028).
@@
The method of equal shares shows stronger legitimacy. The significant legitimacy of equal shares is reflected in several ways (see also Figure LABEL:fig:legitimacy_fairness_preference_heat): (i) Losers, determined as supporters of the top popular losing project (‘Wild bees’ paradise’), prefer the method of equal shares and find it fairer. (ii) Remarkably, the method of equal shares is even found fairer and preferred to a higher extent among losing voters than among winning ones, who often prefer the standard majoritarian voting method. (iii) The shift to equal shares after voting is found fairer among those who do not accept the voting outcome, which is a strong proxy of legitimacy [Weatherford1992, Johnson2006]. However, we also found some cases where more voters shifted to the standard majoritarian voting method after voting (see Section LABEL:sec:choice-ballot-aggregation).
@@
Voting outcome itself vs. method explanation. The explanation of the voting method has an additional contribution of 5.3% to the shift towards equal shares as the preferred method, while it has an additional contribution of 5.2% to the shift towards equal share as the fairer method. Accordingly, the method explanation has a small but statistically significant contribution to the shift to equal shares (with the exception of the mixed setting for the preferred method, p=0.07). These results demonstrate that the voting outcome itself is sufficient to establish a change in favor of the method of equal shares.
@@
2.4 Democratic values explain voters’ shift to the fairer voting method
@@
Clash between democratic values and democratic backsliding. Figure 4 identifies human factors that causally explain the shift from one ballot aggregation method before voting to another one after voting (with the explanation of method), see also Figure 3 and Section 4.3. The classification model is illustrated in Section LABEL:sec:shifts-methods and comes with an accuracy of 85.92%. Strikingly, the group of voters who shift from the standard majoritarian voting method to equal shares exhibits democratic values that explain this transition, see Figure 4c. These voters come with distinguishing qualities (Figure 4b) such as being altruistic, being compromisers, and supporting more expressive, inclusive ballots. They are aware of the proposed projects and flexible to select low-cost projects, which also makes them more likely to oppose a popular costly project. This group is well-represented in the voting outcome. In contrast, the voters that shift from equal shares to the standard majoritarian method do not show these democratic values. Instead they are driven by self-interest, mistrust and dissatisfaction. They support simpler ballot formats, are not well-represented in the voting outcomes, and support the popular loser project under equal shares.
@@
Figure 4: The shift to equal shares as the preferred method after voting is explained by distinguishing human traits that reflect the promotion of strong democratic values such as being altruistic, compromisers, and supporting more expressive, inclusive ballots. In contrast, shifting to the standard majoritarian method is found among voters showing mistrust, under-representation, dissatisfaction, and self-interest. These voters also support the popular costly loser project. A classification model (with 85.92% accuracy) is designed with four classes representing the preferred ballot aggregation method before and after voting (with explanation of methods). The independent variables are human traits of voters that explain the observed transitions. (A) Distinguishing human traits of voters who shift to the standard method and equal shares after voting. These are prominently associated with democratic backsliding on the one hand, and promotion of democratic values on the other hand. (B) Summary of distinguishing qualities for those who prefer the method of equal shares over the standard majoritarian method after voting. Altruism, support of expressive, inclusive ballots and opposition to the popular, costly loser project of equal shares are qualities with the highest prominence for those voters preferring equal shares. (C) Human traits (features) that explain the preference of each voter before and after voting, in particular the transitions from the standard majoritarian method to equal shares and from equal shares to the standard method (see Figure 3). The ‘*’ indicates that a feature is statistically significant (p<0.05) in explaining the voters’ choice of the method before and after voting.
@@
Distinguishing qualities of voters who shift preference to equal share. Figure 4b summarizes the distinguishing qualities of voters, who shift to equal shares by taking the difference of the human traits for the two transitions (standard to equal shares – equal shares to standard). Altruism, opposition to the popular (costly) loser project and support of the expressive, inclusive ballot formats are the most prominent distinguishing qualities observed for voters who shift to equal shares. Figure 4c shows the importance of features that explain the classification of voters based on which of the two transitions they make for their preferred ballot aggregation method before and after voting (see Tables LABEL:tab:accuracy_resultsLABEL:tab:feature-loss and LABEL:tab:trait-survey-map in Section LABEL:sec:shifts-methods). Further results for the transitions to methods that are found fairer and more complex are shown in Figures LABEL:fig:ES_G_G_ESLABEL:fig:ES_GLABEL:fig:Convenience_Explain in Section LABEL:sec:shifts-methods.
@@ 3 Discussion
A blueprint to upgrade democracies. In our study, we have shown how fair voting methods can better foster collective intelligence that is based on higher inclusion and diversity, with strong evidence provided by (i) a rigorous participatory budgeting process in the city of Aarau in Switzerland as well as (ii) extensive comparisons with other participatory processes. Subsequent processes using fair voting methods confirm our findings [equalsharesmethod]. Accordingly, the standard majoritarian approach of voting has a number of deficits, such as the selection of few projects only, which often lack cost-effectiveness, leaving a large portion of good project ideas unrealized. As projects selected by majoritarian voting are typically implemented in the city center, they largely fail to achieve spatial fairness or inclusion. In the past, this issue has been addressed by dividing a city into multiple districts and running multiple budgeting processes in parallel, as in the case of Zurich in 2021. In contrast, combining cumulative voting with the method of equal shares proves to be resourceful, as it supports more winning projects. It also promotes fairness, satisfaction, inclusion, trust, and democratic values. Accordingly, this approach offers a powerful way of upgrading democracies. Implementing this approach, using digital means, is also a value-sensitive approach to promoting digital democracy by design [Helbing2023].
@@
Success of project implementation. By the end of 2024, 9 of 17 projects selected by the “City Idea” process in Aarau were already implemented and 5 projects got an extension, while 3 projects could not be implemented mainly due to feasibility and budgetary constraints [CityIdeaReport2025]. These were among the top-11 projects. Two thirds of these projects would also have been winners if the standard majoritarian method had been used. These results demonstrate that having more winning projects using the method of equal shares does not negatively affect the overall capacity of citizens to implement them. In contrast, the standard majoritarian voting method applied in Zurich for the “City Idea” process did not produce so satisfactory results, and the press critically discussed the problem of cost-inefficient projects [ZurichPBReport].
@@
Participation levels. The participatory budgeting process in Aarau achieved significant participation of citizens at different levels. During the project ideation phase, 161 ideas were proposed, which exceeded by far the initial goal of the campaign to attract 100 ideas [CityIdeaReport2025]. In comparison, 176 ideas were proposed and 135 of them presented for voting in Zurich in 2021 with a 19 times larger population and a 10 times larger total budget. The total of 1,703 participating voters in Aarau (as compared to 1,804 in Zurich) is a significant number for such a campaign in a town of 14,310 residents with voting rights in 2023. Nevertheless, the turnout did not reach the level of official elections such as referenda and national elections, yet. A 22% of voters showed interest in participating in the “City Idea” process even though they were not official residents of Aarau, while 9% of voters were not Swiss citizens. The average age of voters was 43, and there were more female than male voters. Children over 15 years were allowed to vote. Among the low number of citizens who abstained but participated in the surveys (296), the top-5 reasons for not voting were "no time" (17.1%), "forgetting" (12.8%), "lack of information" (9.9%), "unappealing projects" (7.7%) or some "other reason" (7.0%), see Question EB.7 in Table LABEL:table:exit-digital.
@@
Upgrading democracies requires upgrading digital tools. We created and used a novel combination of open-source and value-sensitive digital tools to meet a broad range of expectations regarding transparency, security, privacy, reproducibility, explainability and fairness. This required a significant implementation effort, a mobilization of various communities, a strong commitment and the building of shared values with the city authorities, which is a process that has not been systematically pursued to this extent before. A highlight of this effort was the privacy-preserving verification of citizens’ eligibility when voting with the social security number, which ensured a trustworthy voting procedure. The integration of our open-source implementation of cumulative voting [SPBStadtidee2023] into the Stanford Participatory Budgeting platform is a stepping stone towards participatory tools that are more standardized and interoperable. This is also relevant for other large open-source communities such as Decidim and Consul as well as the ones created by the computational social choice community such as Pabutools [pabutools] and Preflib [Mattei2013].
@@
Explaining and adopting fair voting methods. 53.17% of voters understood how equal shares calculates the winning projects (MB.3.3), while 55.95% understood why popular projects are not easily funded by equal shares (MB.3.6), and 41.26% found this fair (MB.3.7). Strikingly, the results of Figure 3 demonstrate that explaining the expected difference in the voting outcomes reached by equal shares is sufficient for voters to prefer this method. A more detailed explanation of how the method works increases the tendency towards preferring equal shares over the standard method only slightly. This result has significant implications for overcoming the long-standing hesitation with regard to democratic innovations, since citizens show a readiness to accept more complex decision-making methods as long as they understand the anticipated impact on the decision outcome. In addition, the experience of the different results helps citizens to shape a practical understanding of the method. Moreover, given that equal shares is designed for a proportional representation of voter preferences and is known for its stability [Fairstein2023], it is more unlikely to cause controversies over voting outcomes. Nevertheless, the additional effort required to explain the method of equal shares should not be under-estimated. As shown in Section LABEL:sec:equal-shares-explanation, the introduction of equal shares requires a systematic communication effort over the lifetime of the participatory budgeting process, using creative media and visualizations suited for the broader public. At the project ideation phase, awareness of the method is required to mobilize the population for ideas that may not have been considered under the majoritarian rule. As the costs of the proposed projects play a key role for selecting them, citizens need to be resourceful and prioritize different kinds of ideas, which requires a focus on cost-effective projects that are supported by a sufficiently large group of the population. Instead, large and costly projects may be more risky to propose as they require the support of a broader population in order to get selected.